OpenStudy (anonymous):

Simplify x squared minus 12 times x plus 20 all over x minus 10.

OpenStudy (anonymous):

\[\frac{x^2-12x+20}{x-10}\] if you can simplify this at all then it must be the case that you can factor the numerator as \[\frac{(x-10)(\text{something})}{x-10}=\text{something}\] your job is therefore to factor